OpenStudy (anonymous):

an annual interest rate of 18%. Assume that you have a previous charged balance of $285.76 (before interest has been applied) on the card. On your shopping trip, you purchased three items: a Blu-ray player, two 4-GB flash drives, and a 19-inch flat-screen television. You purchase all the items on your credit card for a total of $352.18. When the bill comes at the end of the month you decide to pay all of the charges. What is the monthly interest rate? What are the interest charges on the card? (Calculate only on last month’s unpaid balance) What is the total balance on your card?

OpenStudy (anonymous):

Alright, so lets see....first we need to find out the monthly interest rate. Well, i want to say divide the annual interest rate by 12 (for each month) but that seems a little too nice. Let me read up on something real fast to make sure thats correct (or incorrect)

OpenStudy (anonymous):

Yeah, thats the correct method. So lets do that: \[\frac{.18}{12} = .015\] So thats a monthly interest rate of 1.5%

OpenStudy (anonymous):

Now to figure out what the interest charges on the card. First we need to know the balance on your card. That would be: \[285.76+352.18 =637.94 \] Now lets add on the interest: \[637.94*(.015) = 9.57\] That is just the interest charged.

OpenStudy (anonymous):

Then for the total amount your bill will be, you add the amount of your charges with the interest: \[637.94+9.57 = 647.51\] thats how much the bill will be for.